OpenStudy (anonymous):

How many milligrams of pure water must be added to 6 mg of a 40% solution to obtain a 25% solution? Round to the nearest tenth. 2.8 mg 3.6 mg 3.1 mg 4.3 mg

OpenStudy (anonymous):

3.6 mg

OpenStudy (saifoo.khan):

6x.4=2.4 .25p=2.4 p=2.4/.25=9.6 9.6-6=3.6 ms pure water necessary to be added..

OpenStudy (amistre64):

eliminate some more :) w + 6 = t ;*-.25 w(0) + 6(.4) = t(.25) -w(.25)- 6(.25) = -t(.25) w(0) + 6(.4) = t(.25) ---------------------- -w(.25) +6(.15) = 0 ; solve for w w = 90/25 = 18/5 = 3.6